# PRISM: Leveraging Prototype Patient Representations with Feature-Missing-Aware Calibration for EHR Data Sparsity Mitigation

### v4

@@ Abstract
Electronic Health Record (EHR) data, while rich in information, often suffers from sparsity, posing significant challenges in predictive modeling. Traditional imputation methods inadequately distinguish between real and imputed data, leading to potential inaccuracies in models. Addressing this, we introduce PRISM, a framework that indirectly imputes data through prototype representations of similar patients, thus ensuring denser and more accurate embeddings. PRISM also includes a feature confidence learner module, which evaluates the reliability of each feature in light of missing data. Additionally, it incorporates a new patient similarity metric that accounts for feature confidence, avoiding overreliance on imprecise imputed values. Our extensive experiments on the MIMIC-III, MIMIC-IV, PhysioNet Challenge 2012, eICU datasets demonstrate PRISM’s superior performance in predicting in-hospital mortality and 30-day readmission tasks, showcasing its effectiveness in handling EHR data sparsity. For the sake of reproducibility and further research, we have made the code publicly available at https://github.com/yhzhu99/PRISM.
@@ 1 Introduction
Electronic Health Records (EHR) have become indispensable in modern healthcare, offering a rich source of data that chronicles a patient’s medical history. Over recent years, machine learning techniques have garnered significant attention in leveraging EHR data to inform clinical decisions. Such applications range from predicting the survival risk of patients [1, 2] to forecasting early mortality outcomes [3, 4, 5, 6].
Working with time-series EHR data presents challenges due to its inherent sparsity. Factors such as data corruption [7], expensive examinations [8], and safety considerations [9] result in missing observations; for instance, not all indicators are captured during every patient visit [10]. Imputed values, while necessary, are not genuine reflections of a patient’s condition and can introduce noise, diminishing model accuracy [11]. Given that most machine learning models cannot process NaN (Not a Number) inputs, this sparsity necessitates imputation, complicating EHR predictive modeling. While most existing works tackling EHR sparsity have tried to perform the imputation task directly on raw data, based on modeling the health status trajectory of the whole training set, this approach is similar to that of matrix completion methods, such as MICE [12], non-negative matrix factorization (NMF) [13], and compressed sensing [14]. However, they fail to capture the time-series characteristics of longitudinal EHR data for each patient. Also, previous EHR-specific models with strategies of recalibrating patient representations based on attentive feature importance [15, 16] exhibit an issue of inadvertently prioritizing imputed features, potentially introducing inaccuracies [17]. Thus, addressing the sparsity in time-series EHR data requires a focus on capturing relevant feature representations across visits within a patient or among patients that are essential for predictive purposes. The primary objective is to discern and highlight crucial features while diminishing the impact of irrelevant, redundant, or missing ones.
@@
Given these insights, we are confronted with a pressing challenge: How can we effectively mitigate the sparsity issue in EHR data caused by missing recorded features, ensuring accurate and dense patient representations?
To address this, we introduce PRISM model that leverages prototype similar patients representations at the hidden state level. Unlike traditional direct imputation methods, our strategy, guided by prediction targets and with feature-missing-aware calibration, serves as a more effective imputation strategy, thus mitigating EHR data sparsity issue. Central to our approach is the feature-missing-aware calibration process in our proposed feature confidence learner module. It evaluates the reliability of each feature, considering its absence, the time since the last recorded visit, and the overall rate of missing data in the dataset. By emphasizing feature confidence, our newly designed patient similarity measure provides evaluations based not just on raw data values, but also on the varying confidence levels of each feature. Overall, our indirect imputation method distinctively differs from direct imputation methods, minimizing the introduction of misleading information.
@@ 4.1 Overview
Figure 1 shows the overall pipeline of PRISM. It consists of four main sub-modules below.
@@
  • Confidence-Aware Prototype Patient Learner improves the measure of patient similarity based on patient representation and the confidence level of features learned from feature missing status. It then applies the graph neural network to learn prototype patients.
  • Prototype Representation Fusion Module fuses the prototype patient embedding and the patient’s own representation, further enhancing the hidden state representation of the patient that is affected by missing data.
4.2 Feature-Isolated Embedding
@@
Static Features Embedding Static features remain constant at each visit. Hence, we opt for a single-layer MLP for simplicity to map each static feature into the feature dimensions $f$:
@@
Dynamic Features Embedding To ensure that each feature’s individual statistics, e.g. missing status can be incorporated with the corresponding feature, we adopt multi-channel GRU structure to avoid feature interaction at this stage. Each feature is embed with an isolated GRU, a time-series model that has a proven track record of consistent performance in EHR modeling [5]:
@@ 4.3 Feature Confidence Learner
Feature Missing Status Representation We introduce two feature missing status-related variables $\rho$ and $\tau$. Global missing representation $\rho_{i}$ represents the presence rate of the $i$-th feature in the original dataset. For example, if the dataset collects a total of 100 data records during the visits of all patients, but certain feature is only recorded two times, then the $\rho$ for this feature is $\frac{2}{100}=0.02$. Local missing representation $\tau_{i,t}$ represents the time interval since the last record of this feature at the current visit. There are two special cases: case 1) If the feature is recorded at the current visit, it is marked as 0; case 2) if the feature has never been recorded before the current visit, it is marked as infinity:
@@
Missing-Aware Self-Attention To calculate the feature confidence, we comprehensively consider the missing feature status in the dataset, including the global missing representation $\rho$ and local missing representation $\tau$, and integrate them into a self-attention mechanism module.
@@ 4.4 Confidence-Aware Prototype Patient Learner
To comprehensively account for the implications of absent features in the evaluation and identification of analogous patients, we integrate the feature confidence status into our confidence-aware module for discovering cohorts of similar patients. Inspired by GRASP [2]’s graph convolutional network (GCN) framework in EHR modeling, we further compute and integrate the similarity score within the GCN by factoring in the acquired feature confidence pertaining to the individual patient, aiming to facilitate the preferential selection of patients who are more similar to the subject patient.
Confidence-Aware Patient Similarity Measure To find similar patients for the current patient, we first need to calculate the similarity with other patients. This is achieved through the confidence-aware patient similarity measure, resulting in a similarity score matrix as $\bm{A}$. Specifically, the similarity score $\phi_{i,j}(\bm{z}_{i},\bm{z}_{j};\bm{C}_{i},\bm{C}_{j})$ between the $i$-th and $j$-th ($i\neq j$) patients is defined as Equation 14. Note that when $i=j$, indicating the comparison of a patient with themselves, the similarity score is defined in $\phi_{i,j}=1$.
@@
Prototype Patients Cohort Discovery To compute the enhanced representation information of similar patients, we design the prototype similar patient cohort discovery module by incorporating GCN’s capability to learn relationships between graph nodes. Initially, we utilize the K-Means clustering algorithm to cluster raw patient representation ($F$ recorded or imputed features) into $K$ groups. Subsequently, we identify the center vectors of $K$ clusters as the initial $K$ prototypes $\bm{\mathcal{G}_{k}},k=1,2,\cdots,K$. These prototypes vectors, along with the patient groups, form the nodes of a graph. We then calculate the edge weights for this graph using a predefined similarity measure, resulting in an adjacency matrix $\bm{A}$. As GCN learns across epochs, the graph structure dynamically evolves. Note that the feature confidence $\mathbf{C}$ of prototype patients is initially set to $1$ for each feature. During the training phase, both $\bm{\mathcal{G}}$ and the corresponding $\mathbf{C}$ are adaptively adjusted. Consequently, the most similar and representative prototypes $\bm{\mathcal{G}}$ are identified, and samples within the clusters become more similar to each other. The process is illustrated as:
@@
4.5 Prototype Representation Fusion Module
@@ 5.3 Baseline Models
Imputation-based methods
include MICE [12] addresses missing data in EHR through iterative imputation, with subsequent analysis using an LSTM model [29]. GRU-D [10] incoporates both the last observed and global mean values in the GRU network. Additionally, GRU-D utilizes an exponential decay mechanism to manage the temporal dynamics of missing values.
EHR-specific models
such as RETAIN [30], AdaCare [15], ConCare [1], GRASP [2], M3Care [11], SAFARI [4] and AICare [16]. They focus on personalized health status embeddings.
Ablation models
include PRISM-${}_{\text{-proto.}}$ and PRISM-${}_{\text{-calib.}}$. Specifically, PRISM-${}_{\text{-proto.}}$ removes the confidence-aware prototype similar patient learner and reserves the feature-missing-aware calibration process. We apply the patients hidden representation $\bm{z}$ for downstream GRU module. PRISM-${}_{\text{-calib.}}$ removes the feature confidence learner from the missing-feature-aware calibration process. As the confidence-aware prototype patient learner requires the feature confidence $\bm{C}$ as input, we set the $\bm{C}=1$ for each feature.
@@ 5.4 Implementation Details
All runs are trained on a single Nvidia RTX 3090 GPU with CUDA 11.8. The server’s RAM size is 64GB. We implement the model in Python 3.11.4, PyTorch 2.0.1 [31], PyTorch Lightning 2.0.5 [32], and pyehr [33]. AdamW [34] is employed with a batch size of 1024 patients. All models are trained for 50 epochs with an early stopping strategy based on AUPRC after 10 epochs without improvement. The learning rate ${0.01,0.001,0.0001}$ and hidden dimensions ${64,128}$ are tuned using a grid search strategy on the validation set. The searched hyperparameter for PRISM is: 128 hidden dimensions, 0.001 learning rate, and 256 prototype patient cluster numbers. Performance is reported in the form of mean±std of 5 runs with random seeds ${0,1,2,3,4}$ for MIMIC-III and MIMIC-IV datasets, and apply bootstrapping on all test set samples 10 times for the Challenge-2012 and eICU datasets, following practices in [16]. All metric scores are multiplied by 100 for readability purposes.
@@ 6 Experimental Results and Analysis
We conduct the in-hospital mortality and 30-day readmission prediction task on MIMIC-III and MIMIC-IV datasets, in-hospital mortality prediction task on Challenge-2012 and eICU datasets. The dataset preprocessing scripts, developed model’s code and all baseline models’ code are open-sourced for reproducibility.
Table 3: Benchmarking results on MIMIC-III, MIMIC-IV, Challenge-2012, and eICU datasets. Bold indicates the best performance.
@@ 6.2 Ablation Study
Comparing with Reduced Versions
@@
Comparing with Internal Components
@@
Table 4: Performance comparison of internal components on the MIMIC-IV mortality prediction task. “Feat. Conf.” means “Feature Confidence” and “Sim. Meas.” denotes “Similarity Measure”. Bold denotes the best performance within each components, Red denotes the highest performance among all comparisons.
@@ 6.3 Observations and Analysis
Robustness to Data Sparsity
@@
Sensitiveness to Cohort Size and Effectiveness of Missing-Feature-Aware Module
@@
Variations of Similarity Measures
We conduct a comparison against standard similarity metrics, as detailed in Table 5, PRISM outperforms cosine similarity, and L1 and L2 distances in all metrics. This demonstrates the effectiveness of PRISM’s confidence-aware patient similarity measure in handling sparse EHR data.
Table 5: Comparing different similarity measures in MIMIC-IV in-hospital mortality prediction task.
@@
Model Efficiency and Complexity
We evaluate the efficiency and complexity against SOTA EHR-specific models baseline models. Table 6 highlights PRISM’s balance between a relatively low total parameter count (215K) and a competitive runtime (69.03s).

### v5

@@ Abstract
Electronic Health Records (EHRs) contain a wealth of patient data; however, the sparsity of EHRs data often presents significant challenges for predictive modeling. Conventional imputation methods inadequately distinguish between real and imputed data, leading to potential inaccuracies of patient representations. To address these issues, we introduce PRISM, a framework that indirectly imputes data by leveraging prototype representations of similar patients, thus ensuring compact representations that preserve patient information. PRISM also includes a feature confidence learner module, which evaluates the reliability of each feature considering missing statuses. Additionally, PRISM introduces a new patient similarity metric that accounts for feature confidence, avoiding overreliance on imprecise imputed values. Our extensive experiments on the MIMIC-III, MIMIC-IV, PhysioNet Challenge 2012, eICU datasets demonstrate PRISM’s superior performance in predicting in-hospital mortality and 30-day readmission tasks, showcasing its effectiveness in handling EHR data sparsity. For the sake of reproducibility and further research, we have made the code publicly available at https://github.com/yhzhu99/PRISM.
@@ 1 Introduction
Electronic Health Records (EHR) have become indispensable in modern healthcare, offering a rich source of data that chronicles a patient’s medical history. Over recent years, machine learning techniques have gained significant attention for their ability to leverage time-series EHR data, which represented as temporal sequences of high-dimensional clinical variables [1], can significantly inform and enhance clinical decision-making. Such applications range from predicting the survival risk of patients [2, 3] to forecasting early mortality outcomes [4, 5, 6, 7].
Working with time-series EHR data presents challenges due to its inherent sparsity. Factors such as data corruption [8], expensive examinations [9], and safety considerations [10] result in missing observations; for instance, not all indicators are captured during every patient visit [11]. Imputed values, while necessary, are not genuine reflections of a patient’s condition and can introduce noise, diminishing model accuracy [12]. Given that most machine learning models cannot process NaN (Not a Number) inputs, this sparsity necessitates imputation, complicating EHR predictive modeling. While most existing works tackling EHR sparsity have tried to perform the imputation task directly on raw data, based on modeling the health status trajectory of the whole training set, this approach is similar to that of matrix completion methods, such as MICE [13], non-negative matrix factorization (NMF) [14], and compressed sensing [15]. However, they fail to capture the temporal interactions of longitudinal EHR data for each patient. Also, previous EHR-specific models with strategies of recalibrating patient representations based on attentive feature importance [16, 17] exhibit an issue of inadvertently prioritizing imputed features, potentially introducing inaccuracies [18]. Thus, addressing the sparsity in time-series EHR data requires a focus on capturing relevant feature representations across visits within a patient or among patients that are essential for predictive purposes. The primary objective is to discern and highlight crucial features while diminishing the impact of irrelevant, redundant, or missing ones.
@@
Given these insights, we are confronted with a pressing challenge: How can we effectively mitigate the sparsity issue in EHR data caused by missing recorded features while ensuring a compact patient representation that preserves patient information? [22]
To address this, we introduce PRISM that leverages prototype similar patients representations at the hidden state space. Unlike traditional direct imputation methods that imputes values based only raw data, PRISM learns refined patient representations according to prediction targets, serving as a more effective imputation strategy and thus mitigating the EHR data sparsity issue. Central to our approach is the feature-missing-aware calibration process in the proposed feature confidence learner module. It evaluates the reliability of each feature, considering its absence, the time since the last recorded visit, and the overall rate of missing data in the dataset. By emphasizing feature confidence, our newly designed patient similarity measure provides evaluations based not just on raw data values, but also on the varying confidence levels of each feature.
@@ 3 Problem Formulation
3.1 EHR Datasets Formulation
3.2 Predictive Objective Formulation
3.3 Notation Table
@@ 4.1 Overview
Figure 1 shows the overall pipeline of PRISM. It consists of three main sub-modules below.
@@
  • Confidence-Aware Prototype Patient Learner improves the measure of patient similarity based on patient representation and the confidence level of features learned from feature missing status. It then applies the graph neural network to learn prototype patients. Finally, the patient’s own representation is fused with the prototype patient representation adaptively, further enhancing the hidden state representation of the patient that is affected by missing data. A two-layer GRU network is then utilized to generate a compact patient health representation, followed by a single-layer MLP network to conduct downstream specific prediction tasks.
4.2 Feature-Isolated Embedding Module
@@
4.2.1 Static Features Embedding
Static features remain constant at each visit. Hence, we opt for a single-layer MLP for simplicity to map each static feature into the feature dimensions $f$:
@@ 4.2.1 Static Features Embedding
4.2.2 Dynamic Features Embedding
To ensure that each feature’s individual statistics, e.g. missing status can be incorporated with the corresponding feature, we adopt multi-channel GRU structure to avoid feature interaction at this stage. Each feature is embed with an isolated GRU, a time-series model that has a proven track record of consistent performance in EHR modeling [6]:
@@ 4.3 Feature Confidence Learner
4.3.1 Feature Missing Status Representation
We introduce $\rho$ and $\tau$ to record feature missing status. Global missing representation $\rho_{i}$ represents the presence rate of the $i$-th feature in the original dataset:
$\rho_{i}=\frac{\text{total observations of}\ i\text{-th feature}}{\text{total visits of all patients}}$ (4)
For example, if the dataset collects a total of 100 data records during the visits of all patients, but certain feature is only recorded two times, then the $\rho$ for this feature is $\frac{2}{100}=0.02$. Local missing representation $\tau_{i,t}$ represents the time interval since the last record of this feature at the current visit. There are two special cases: case 1) If the feature is recorded at the current visit, it is marked as 0; case 2) if the feature has never been recorded before the current visit, it is marked as infinity:
@@ 4.3.1 Feature Missing Status Representation
4.3.2 Missing-Aware Self-Attention
To calculate the feature confidence, we comprehensively consider the missing feature status in the dataset, including the global missing representation $\rho$ and local missing representation $\tau$, and integrate them into a self-attention mechanism module.
@@ 4.4 Confidence-Aware Prototype Patient Learner
We incorporate feature confidence status into our confidence-aware module for identifying similar patient cohorts, accounting for the impact of missing features. Inspired by GRASP’s [3] graph convolutional network (GCN) framework, we further compute and integrate the similarity score within the GCN, considering the individual patient’s feature confidence to prioritize the selection of patients most similar to the subject patient.
4.4.1 Confidence-Aware Patient Similarity Measure
To find similar patients, we calculate the similarity between the current patient and others using the confidence-aware patient similarity measure, resulting in a similarity score matrix $\bm{A}$. Specifically, the similarity score $\phi_{i,j}(\bm{z}_{i},\bm{z}_{j};\bm{C}_{i},\bm{C}_{j})$ between the $i$-th and $j$-th ($i\neq j$) patients is defined as Equation 15. Note that when $i=j$, indicating the comparison of a patient with themselves, the similarity score is defined in $\phi_{i,j}=1$.
@@ 4.4.1 Confidence-Aware Patient Similarity Measure
4.4.2 Prototype Patients Cohort Discovery
To compute the enhanced representations of similar patients, we design the prototype patients cohort discovery module by incorporating GCN’s capability to learn relationships between graph nodes. Initially, we utilize the K-Means clustering algorithm to cluster raw patient representation ($F$ recorded or imputed features) into $K$ groups. Subsequently, we identify the center vectors of $K$ clusters as the initial $K$ prototypes $\bm{\mathcal{G}_{k}},k=1,2,\cdots,K$. These prototypes vectors, along with the patient groups, form the nodes of a graph. We then calculate the edge weights for this graph using a predefined similarity measure, resulting in an adjacency matrix $\bm{A}$. As GCN learns across epochs, the graph structure dynamically evolves. Note that the feature confidence $\mathbf{C}$ of prototype patients is initially set to $1$ for each feature. During the training phase, both $\bm{\mathcal{G}}$ and the corresponding $\mathbf{C}$ are adaptively adjusted. Consequently, the most similar and representative prototypes $\bm{\mathcal{G}}$ are identified, and samples within the clusters become more similar to each other. The process is illustrated as:
@@ 4.4.2 Prototype Patients Cohort Discovery
4.4.3 Prototype Representation Fusion
@@ 5.3 Baseline Models
5.3.1 Imputation-based Methods
We include two imputation-based methods: MICE and GRU-D:
  • MICE [13] addresses missing data in EHR through iterative imputation, with subsequent analysis using an LSTM model [31].
  • GRU-D [11] incoporates both the last observed and global mean values in the GRU network. Additionally, GRU-D utilizes an exponential decay mechanism to manage the temporal dynamics of missing values.
5.3.2 EHR-specific Models
Following methods are specifically designed for EHR data and focus on personalized health status embeddings.
  • RETAIN [1] is a hierarchical attention-based interpretable model. It attends the EHR data in a reverse time order so that recent clinical visits are likely to receive higher attention.
  • AdaCare [16] is a GRU-based network that utilizes a multi-scale dilated convolutional module to capture the long and short-term historical variation.
  • ConCare [2] utilizes multi-channel GRU with a time-aware attention mechanism to extract clinical features and re-encode the clinical information by capturing the interdependencies between features.
  • GRASP [3] is a generic framework for healthcare models, which leverages the information extracted from patients with similar conditions to enhance the cohort representation learning results.
  • M3Care [12] resolves the missing modality issue in EHR data by utilizing similar patients’ existing modalities. However, it does not address the issue of missing features within available modalities.
  • SAFARI [5] learns patient health representations by applying a clinical-fact-inspired, task-agnostic correlational sparsity prior to medical feature correlations, using a bi-level optimization process that involves both inter- and intra-group correlations.
  • AICare [17] also includes a multi-channel feature extraction module and an adaptive feature importance recalibration module. It learns personalized health status embeddings with static and dynamic features.
5.3.3 Ablation Models
Ablation models include PRISM-${}_{\text{-proto.}}$ and PRISM-${}_{\text{-calib.}}$.
  • PRISM-${}_{\text{-proto.}}$ removes the confidence-aware prototype similar patient learner and reserves the feature-missing-aware calibration process. We apply the patients hidden representation $\bm{z}$ for downstream GRU module.
  • PRISM-${}_{\text{-calib.}}$ removes the feature confidence learner from the missing-feature-aware calibration process. As the confidence-aware prototype patient learner requires the feature confidence $\bm{C}$ as input, we set the $\bm{C}=1$ for each feature.
@@ 5.4 Implementation Details
5.4.1 Hardware and Software Configuration
All runs are trained on a single Nvidia RTX 3090 GPU with CUDA 11.8. The server’s system memory (RAM) size is 64GB. We implement the model in Python 3.11.4, PyTorch 2.0.1 [32], PyTorch Lightning 2.0.5 [33], and pyehr [34].
5.4.2 Model Training and Hyperparameters
AdamW [35] is employed with a batch size of 1024 patients. All models are trained for 50 epochs with an early stopping strategy based on AUPRC after 10 epochs without improvement. The learning rate ${0.01,0.001,0.0001}$ and hidden dimensions ${64,128}$ are tuned using a grid search strategy on the validation set. The searched hyperparameter for PRISM is: 128 hidden dimensions, 0.001 learning rate, and 256 prototype patient cluster numbers. Performance is reported in the form of mean±std of 5 runs with random seeds ${0,1,2,3,4}$ for MIMIC-III and MIMIC-IV datasets, and apply bootstrapping on all test set samples 10 times for the Challenge-2012 and eICU datasets, following practices in AICare [17].
@@ 6 Experimental Results and Analysis
We conduct the in-hospital mortality and 30-day readmission prediction task on MIMIC-III and MIMIC-IV datasets, in-hospital mortality prediction task on Challenge-2012 and eICU datasets. The dataset preprocessing scripts, developed model’s code and all baseline models’ code are attached at the Supplementary Material for reproducibility.
6.1 Experimental Results
Table 3: Benchmarking results on MIMIC-III, MIMIC-IV, Challenge-2012, and eICU datasets. Bold indicates the best performance. Performance is reported in the form of mean±std. All metric scores are multiplied by 100 for readability purposes.
@@ 6.2 Ablation Study
6.2.1 Comparing with Reduced Versions
@@
6.2.2 Comparing with Internal Components
@@
Table 4: Performance comparison of internal components on the MIMIC-IV mortality prediction task. “Feat. Conf.” means “Feature Confidence” and “Sim. Meas.” denotes “Similarity Measure”. Bold denotes the best performance within each components, Red denotes the highest performance among all comparisons. Performance is reported in the form of mean±std. All metric scores are multiplied by 100 for readability purposes.
@@ 6.3 Observations and Analysis
6.3.1 Robustness to Data Sparsity
@@
6.3.2 Sensitiveness to Cohort Size and Effectiveness of Missing-Feature-Aware Module
@@
6.3.3 Variations of Similarity Measures
We assess the performance of PRISM by comparing it against standard similarity metrics commonly used in evaluating patient similarities. Table 5 details the results of this comparison on the MIMIC-IV in-hospital mortality prediction task. As demonstrated in the table, PRISM, leveraging a confidence-aware patient similarity measure, consistently surpasses traditional metrics such as cosine similarity, and L1 and L2 distances. This showcases the effectiveness of PRISM’s similarity metric in measuring the impact of missing features.
Table 5: Comparing different similarity measures in MIMIC-IV in-hospital mortality prediction task. Performance is reported in the form of mean±std. All metric scores are multiplied by 100 for readability purposes.
@@ 6.3.3 Variations of Similarity Measures
6.3.4 Model Efficiency and Complexity
We evaluate the efficiency and complexity of PRISM against SOTA EHR-specific models baseline models. Table 6 compares the total parameter count and runtime of PRISM with these models on the MIMIC-IV dataset, using the same hidden dimension (128) and batch size (1024) hyperparameters. PRISM achieves a competitive balance between a low parameter count (215K) and an efficient runtime (69.03s).